# Finite-Time Analysis of Gradient Descent for Shallow Transformers

### v1

@@ A.6 Proof of Lemma 5
Fix $X\in\mathcal{X}$ and write $a(X;W)=:a$, $M(X;W)=:M$. The linear model evaluated at $\tilde{\varphi}$ decomposes as
$f_{\mathrm{lin}}(X;\tilde{\varphi})\;=\;\frac{1}{m}\sum_{i=1}^{m}g_{i}(X),$
where, with $\langle\cdot,\cdot\rangle_{F}$ the Frobenius product,
$g_{i}(X)$ $:=\phi_{c}\big(X;\varphi^{(0)}_{i}\big)\,v_{c}\!\big(\varphi^{(0)}_{i}\big)+\big\langle\phi_{u}\big(X;\varphi^{(0)}_{i}\big),\,v_{u}\!\big(\varphi^{(0)}_{i}\big)\big\rangle+\big\langle\phi_{w}\big(X;\varphi^{(0)}_{i}\big),\,v_{w}\!\big(\varphi^{(0)}_{i}\big)\big\rangle_{F},$ (19)
$\phi_{c}(X;\varphi_{0})$ $:=\sigma\!\big(U^{\top}a(X;W)\big),$
$\phi_{u}(X;\varphi_{0})$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,a(X;W),$
$\phi_{w}(X;\varphi_{0})$ $:=\sigma^{\prime}\!\big(U^{\top}a(X;W)\big)\,(M(X;W)\,U)\,q_{X}^{\top}.$
By the definition of $\tilde{f}(\cdot;v)$,
$\tilde{f}(X;v)\;:=\;\mathbb{E}\!\big[\phi_{c}(X;\varphi_{0})\,v_{c}(\varphi_{0})\big]+\mathbb{E}\!\big[\langle\phi_{u}(X;\varphi_{0}),v_{u}(\varphi_{0})\rangle\big]+\mathbb{E}\!\big[\langle\phi_{w}(X;\varphi_{0}),v_{w}(\varphi_{0})\rangle_{F}\big]\;=\;\mathbb{E}\!\big[g_{1}(X)\big],$ (20)
where the expectation is over a single draw $\varphi_{0}\sim\varphi^{(0)}_{1}$. Hence
$f_{\mathrm{lin}}(X;\tilde{\varphi})-\tilde{f}(X;v)\;=\;\frac{1}{m}\sum_{i=1}^{m}\Big(g_{i}(X)-\mathbb{E}[g_{i}(X)]\Big).$ (21)
Recall $m$ is even and $(\varphi^{(0)}_{i})_{i\leq m/2}$ are i.i.d., while $\varphi^{(0)}_{i+m/2}$ equals $\varphi^{(0)}_{i}$ with the $c_{i}$ sign-flipped. Define
$\bar{g}_{i}(X)\;:=\;\frac{g_{i}(X)+g_{i+m/2}(X)}{2},\qquad i=1,\dots,m/2.$
Then $\{\bar{g}_{i}(X)\}_{i=1}^{m/2}$ are i.i.d. and $\mathbb{E}\bar{g}_{1}(X)=\mathbb{E}g_{1}(X)$, so
$\Bigg|\frac{1}{m}\sum_{i=1}^{m}g_{i}(X)-\mathbb{E}[g_{1}(X)]\Bigg|\;=\;\Bigg|\,\frac{1}{m/2}\sum_{i=1}^{m/2}\bar{g}_{i}(X)-\mathbb{E}[\bar{g}_{1}(X)]\,\Bigg|.$ (22)
We use the Orlicz $\psi_{2}$ norm (see, e.g., (Vershynin, 2025, Chap. 2)):
$\|Z\|_{\psi_{2}}\;:=\;\inf\{t>0:\ \mathbb{E}\exp(Z^{2}/t^{2})\leq 2\}.$
(i) $c$ term. Since $|v_{c}(\varphi_{0})\phi_{c}|\leq\sigma_{0}\bar{\nu}_{c}$ a.s., then $\|\phi_{c}\,v_{c}\|_{\psi_{2}}\;\leq\;\sigma_{0}\,\bar{\nu}_{c}.$
(ii) $U$ term. Since $|\langle\phi_{u},v_{u}\rangle|\leq\|\phi_{u}\|_{2}\|v_{u}\|_{2}\leq\sigma_{1}\bar{\nu}_{u}$ a.s., we have $\|\langle\phi_{u},v_{u}\rangle\|_{\psi_{2}}\;\leq\;\sigma_{1}\,\bar{\nu}_{u}.$
(iii) $W$ term. We can write $\langle\phi_{w},v_{w}\rangle_{F}=\langle U,\sigma^{\prime}(U^{\top}a)M\,v_{w}(\varphi_{0})\,q_{X}\rangle$. Using the fact that $U$ is Gaussian and $\|\sigma^{\prime}(U^{\top}a)M\,v_{w}(\varphi_{0})\,q_{X}\|_{2}\leq\bar{\nu}_{w}$ a.s., we have $\|\langle\phi_{w},v_{w}\rangle_{F}\|_{\psi_{2}}\;\leq\;\sigma_{1}\;\sqrt{\tfrac{8}{3}}\;\bar{\nu}_{w}\;\leq\;2\sigma_{1}\bar{\nu}_{w}.$
Combining (i)–(iii) by the triangle inequality for $\psi_{2}$ norms,
$\|g_{i}(X)\|_{\psi_{2}}\;\leq\;K\quad\text{with}\quad K:=2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w}).$ (23)
Since $\{\bar{g}_{i}(X)\}_{i=1}^{m/2}$ are i.i.d. and centered after subtracting $\mathbb{E}\bar{g}_{1}(X)$, the standard sub-Gaussian tail bound yields, for all $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{m/2}\sum_{i=1}^{m/2}\big(\bar{g}_{i}(X)-\mathbb{E}\bar{g}_{1}(X)\big)\right|\geq t\right)\;\leq\;2\exp\!\left(-\,\frac{(m/2)\,t^{2}}{K^{2}}\right).$
@@
Choosing $t\;=\;K\,\sqrt{\tfrac{2\log(2n/\delta)}{m}}$ and union bounding over $X^{(j)}$, we obtain with probability at least $1-\delta$,
@@
$\sup_{j\in[n]}\Big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\Big|\;\leq\;2\Big(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w}\Big)\,\sqrt{\frac{2\log(2n/\delta)}{m}}.$
@@ A.7 Proof of Theorem 1
Set the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$ and write the pre-projection point $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})$. By non-expansivity of Euclidean projection onto a convex set and the Pythagorean identity,
@@ A.8 Proof of Proposition 1
Under the same conditions as Theorem 1, we have
@@
Define the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$ and the pre‑projection iterate $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})$. By non‑expansivity of projection and the Pythagorean identity,

### v2

@@ A.6 Proof of Lemma 5
Fix $X\in\mathcal{X}$, and write $a:=a(X;W)$, $M:=M(X;W)$. Define, for each $i\in[m]$,
$g_{i}(X):=g_{i}^{(c)}(X)+g_{i}^{(u)}(X)+g_{i}^{(w)}(X),$
where
$g_{i}^{(c)}(X)$ $:=\phi_{c}\!\big(X;\varphi_{i}^{(0)}\big)\,v_{c}\!\big(\varphi_{i}^{(0)}\big),$
$g_{i}^{(u)}(X)$ $:=\Big\langle\phi_{u}\!\big(X;\varphi_{i}^{(0)}\big),\,v_{u}\!\big(\varphi_{i}^{(0)}\big)\Big\rangle,$
$g_{i}^{(w)}(X)$ $:=\Big\langle\phi_{w}\!\big(X;\varphi_{i}^{(0)}\big),\,v_{w}\!\big(\varphi_{i}^{(0)}\big)\Big\rangle_{F},$
with
$\phi_{c}(X;\varphi_{0})=\sigma(U^{\top}a(X;W)),\qquad\phi_{u}(X;\varphi_{0})=\sigma^{\prime}(U^{\top}a(X;W))\,a(X;W),$
and
$\phi_{w}(X;\varphi_{0})=\sigma^{\prime}(U^{\top}a(X;W))\,(M(X;W)U)\,q_{X}^{\top}.$
By the definition of $f_{\mathrm{lin}}$ and $\tilde{f}(\cdot;v)$,
$f_{\mathrm{lin}}(X;\tilde{\varphi})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(X),\qquad\tilde{f}(X;v)=\mathbb{E}[g_{1}(X)].$
Hence
$f_{\mathrm{lin}}(X;\tilde{\varphi})-\tilde{f}(X;v)=\frac{1}{m}\sum_{i=1}^{m}\big(g_{i}(X)-\mathbb{E}[g_{i}(X)]\big).$
Using the paired initialization, define
$\bar{g}_{i}^{(\ell)}(X):=\frac{1}{2}\Big(g_{i}^{(\ell)}(X)+g_{i+m/2}^{(\ell)}(X)\Big),\qquad\ell\in\{c,u,w\},\quad i\in[m/2].$
Since the pairs $\big(\varphi_{i}^{(0)},\varphi_{i+m/2}^{(0)}\big)$ are i.i.d. across $i\in[m/2]$, the random variables
$Z_{i}(X):=\sum_{\ell\in\{c,u,w\}}\Big(\bar{g}_{i}^{(\ell)}(X)-\mathbb{E}[\bar{g}_{1}^{(\ell)}(X)]\Big),\qquad i\in[m/2],$
are i.i.d. and mean zero, and
$f_{\mathrm{lin}}(X;\tilde{\varphi})-\tilde{f}(X;v)=\frac{2}{m}\sum_{i=1}^{m/2}Z_{i}(X).$
$c$-component.
Since $|\phi_{c}|\leq\sigma_{0}$ and $|v_{c}|\leq\bar{\nu}_{c}$, we have
$|g_{i}^{(c)}(X)|\leq\sigma_{0}\bar{\nu}_{c},\qquad|\bar{g}_{i}^{(c)}(X)|\leq\sigma_{0}\bar{\nu}_{c}.$
Therefore, by Hoeffding’s lemma,
$\mathbb{E}\exp\!\Big(\lambda\big(\bar{g}_{i}^{(c)}(X)-\mathbb{E}\bar{g}_{i}^{(c)}(X)\big)\Big)\leq\exp\!\Big(\frac{\lambda^{2}(\sigma_{0}\bar{\nu}_{c})^{2}}{2}\Big),\qquad\forall\lambda\in\mathbb{R}.$
$u$-component.
By Fact 2, $\|a(X;W)\|_{2}\leq 1$. Since $|\sigma^{\prime}|\leq\sigma_{1}$ and $\|v_{u}\|_{2}\leq\bar{\nu}_{u}$,
$|g_{i}^{(u)}(X)|\leq\|\phi_{u}(X;\varphi_{i}^{(0)})\|_{2}\,\|v_{u}(\varphi_{i}^{(0)})\|_{2}\leq\sigma_{1}\bar{\nu}_{u},$
and hence $|\bar{g}_{i}^{(u)}(X)|\leq\sigma_{1}\bar{\nu}_{u}$. Another application of Hoeffding’s lemma gives
$\mathbb{E}\exp\!\Big(\lambda\big(\bar{g}_{i}^{(u)}(X)-\mathbb{E}\bar{g}_{i}^{(u)}(X)\big)\Big)\leq\exp\!\Big(\frac{\lambda^{2}(\sigma_{1}\bar{\nu}_{u})^{2}}{2}\Big),\qquad\forall\lambda\in\mathbb{R}.$
$w$-component.
Let $Y_{i}:=\bar{g}_{i}^{(w)}(X)$. We have $|Y_{i}|\leq\sigma_{1}\bar{\nu}_{w}\|M(X;W_{i}^{(0)})U_{i}^{(0)}\|_{2}$. We claim that
$\mathbb{E}\exp\!\Big(\lambda(Y_{i}-\mathbb{E}Y_{i})\Big)\leq\exp(2\lambda^{2}(\sigma_{1}\bar{\nu}_{w})^{2}),\qquad\forall\,\lambda\in\mathbb{R}.$
By standard symmetrization trick, it suffices to show
$\mathbb{E}|Y_{i}|^{2k}\leq(\sigma_{1}\bar{\nu}_{w})^{2k}\cdot\frac{(2k)!}{2^{k}k!},\qquad k\geq 1.$ (20)
Conditional on $W_{i}^{(0)}$, let $s_{1},\dots,s_{d}$ be the singular values of $M(X;W_{i}^{(0)})$. Since $U_{i}^{(0)}\sim\mathcal{N}(0,I_{d})$, we have $\|MU_{i}^{(0)}\|_{2}^{2}\stackrel{{\scriptstyle d}}{{=}}\sum_{r=1}^{d}s_{r}^{2}G_{r}^{2}$ with $G_{r}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. By Jensen’s inequality (viewing $\{s_{r}^{2}/\sum s_{r}^{2}\}$ as weights and using convexity of $x\mapsto x^{k}$),
$\Big(\sum_{r}s_{r}^{2}G_{r}^{2}\Big)^{k}\leq\Big(\sum_{r}s_{r}^{2}\Big)^{k-1}\sum_{r}s_{r}^{2}G_{r}^{2k}.$
Taking conditional expectation and using $\sum_{r}s_{r}^{2}=\mathrm{tr}(M^{\top}M)\leq(\mathrm{tr}\,M)^{2}\leq 1$ (Fact 3), we get
$\mathbb{E}\big[\|MU_{i}^{(0)}\|_{2}^{2k}\mid W_{i}^{(0)}\big]\leq\mathbb{E}|G_{1}|^{2k}=\frac{(2k)!}{2^{k}k!},$
which gives (20) after multiplying by $(\sigma_{1}\bar{\nu}_{w})^{2k}$ and taking full expectation.
Substituting (20) into the symmetrized MGF series yields
$\mathbb{E}\,e^{\lambda(Y_{i}-\mathbb{E}Y_{i})}\leq\sum_{k=0}^{\infty}\frac{(2\lambda\sigma_{1}\bar{\nu}_{w})^{2k}}{(2k)!}\cdot\frac{(2k)!}{2^{k}k!}=\sum_{k=0}^{\infty}\frac{(2\lambda^{2}(\sigma_{1}\bar{\nu}_{w})^{2})^{k}}{k!}=e^{2\lambda^{2}(\sigma_{1}\bar{\nu}_{w})^{2}},$
proving the claim.
Combining the bounds.
Write
$Z_{i}^{(c)}:=\bar{g}_{i}^{(c)}-\mathbb{E}\bar{g}_{i}^{(c)},\qquad Z_{i}^{(u)}:=\bar{g}_{i}^{(u)}-\mathbb{E}\bar{g}_{i}^{(u)},\qquad Z_{i}^{(w)}:=\bar{g}_{i}^{(w)}-\mathbb{E}\bar{g}_{i}^{(w)}.$
Then $Z_{i}=Z_{i}^{(c)}+Z_{i}^{(u)}+Z_{i}^{(w)}$. For any $\lambda\in\mathbb{R}$, applying Hölder’s inequality with exponents $1/p_{c}+1/p_{u}+1/p_{w}=1$ we get
$\mathbb{E}e^{\lambda Z_{i}}$ $\leq\Big(\mathbb{E}e^{p_{c}\lambda Z_{i}^{(c)}}\Big)^{1/p_{c}}\Big(\mathbb{E}e^{p_{u}\lambda Z_{i}^{(u)}}\Big)^{1/p_{u}}\Big(\mathbb{E}e^{p_{w}\lambda Z_{i}^{(w)}}\Big)^{1/p_{w}}$
$\leq\exp\!\Big(\frac{p_{c}\lambda^{2}(\sigma_{0}\bar{\nu}_{c})^{2}}{2}\Big)\exp\!\Big(\frac{p_{u}\lambda^{2}(\sigma_{1}\bar{\nu}_{u})^{2}}{2}\Big)\exp\!\Big(2p_{w}\lambda^{2}(\sigma_{1}\bar{\nu}_{w})^{2}\Big)$
$=\exp\!\Big(\frac{\lambda^{2}}{2}(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w})^{2}\Big),$
where in the last equality we set
$p_{c}=\frac{\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w}}{\sigma_{0}\bar{\nu}_{c}},\qquad p_{u}=\frac{\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w}}{\sigma_{1}\bar{\nu}_{u}},\qquad p_{w}=\frac{\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w}}{2\sigma_{1}\bar{\nu}_{w}}.$
Thus each $Z_{i}$ is centered sub-Gaussian with proxy $\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w}$. Since $Z_{1},\dots,Z_{N}$ are i.i.d., for every $\lambda\in\mathbb{R}$,
$\mathbb{E}\exp\!\Big(\lambda\sum_{i=1}^{m/2}Z_{i}\Big)\leq\exp\!\Big(\frac{m\lambda^{2}}{4}(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w})^{2}\Big).$
By Chernoff’s bound,
$\mathbb{P}\!\left(\left|\frac{2}{m}\sum_{i=1}^{m/2}Z_{i}\right|\geq t\right)\leq 2\exp\!\left(-\frac{mt^{2}}{4(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w})^{2}}\right).$
Choosing
$t=\sqrt{2}\,(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w})\sqrt{\frac{2\log(2n/\delta)}{m}},$
@@
and a union bound over $X^{(1)},\dots,X^{(n)}$ yields, with probability at least $1-\delta$,
@@
$\varepsilon_{\mathrm{app}}=\sup_{j\in[n]}\left|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\right|\leq\sqrt{2}\,(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+2\sigma_{1}\bar{\nu}_{w})\sqrt{\frac{2\log(2n/\delta)}{m}}.$
This proves the lemma. ∎
@@ A.7 Proof of Theorem 1
Choose $v$ with $f^{\star}=\tilde{f}(\cdot;v)$ and set the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$, where $\tilde{\varphi}$ is from (11). Write the pre-projection point $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla\widehat{\mathcal{L}}_{n}(\varphi^{(s)})$. By non-expansivity of Euclidean projection onto a convex set and the Pythagorean identity,
@@ A.8 Proof of Proposition 1
Under the same conditions as Theorem 1, we have, on the event $\mathcal{E}_{U,m}(\delta^{\prime})\cap\mathcal{E}_{\mathrm{app}}(\delta)$,
@@
Choose $v$ with $f^{\star}=\tilde{f}(\cdot;v)$ and define the Lyapunov function $\mathcal{V}(\varphi)=\|\varphi-\tilde{\varphi}\|_{2}^{2}$, where $\tilde{\varphi}$ is from (11). Write the pre‑projection iterate $\varphi^{\prime(s)}=\varphi^{(s)}-\eta\,\nabla_{\varphi}\ell_{J_{s}}(\varphi^{(s)})$. By non‑expansivity of projection and the Pythagorean identity,